# High-Precision Physics Experiments at Huizhou Large-Scale Scientific Facilities

### v1

@@ III.3 Probing and Knocking with Muons
Historically, China lacked muon-focused experiments due to limited accelerator capabilities. The High Energy Fragment Separator beamline at HIAF will change this, delivering GeV-scale muon beams [14]. As shown in Fig. 4, this enables studies on muon-philic DM [89, 90], charged lepton flavor violation(CLFV) [91], and quantum correlations in $\mu-e$ scattering [92, 93].
@@
Our experiment will employ Gas Electron Multiplier (GEM) [95] and Resistive Plate Chamber (RPC) [96] detectors for muon tracking. GEMs provide precise spatial resolution, widely applied in high-energy physics, including CMS [97, 98]. RPCs offer timing precision, robustness, and stability. Together, they offer a cost-effective platform for high-resolution tracking.
Dark Matter: We propose a model-independent DM search using cosmic muons tracked by GEMs and RPCs. Simulations with Geant4 [99] suggest muon trajectory shifts upon DM scattering. The CRY program [100] provides a 3–4 GeV average sea-level muon spectrum. Assuming a local DM density of $\rho\sim 0.3$ GeV/$\rm{cm}^{3}$ and a detector volume of $V\sim 1\ {\rm m}^{3}$, the estimated number of DM particles is $3\times 10^{5(6,4)}$. For a one-year run, the expected sensitivity of the muon-philic dark matter scattering cross section is:
@@
This approach complements nuclear recoil searches(e.g., XENON1T, PandaX) and may reveal strongly interacting DM accumulating inside Earth [90]. Detection efficiency remains key to improving limits. To further enhance the sensitivity of DM searches, we can conduct experiments untilizing muon beams with higher intensity and better focusing characteristic.
Light Dark Bosons: Thermal freeze-out models for sub-GeV DM require light mediators, such as a dark $Z^{\prime}$ gauge boson [101]. The $L_{\mu}-L_{\tau}$ model [102, 103, 104, 105], predicts $Z^{\prime}$ coupling weakly to muons. We target the process $\mu e^{-}\to\mu e^{-}X$($X$ decays invisibly), characterized by large-angle muon and electron scattering with minimal downstream signals. The 1–10 GeV muon beam at HIAF-HFRS [14](operational20252026similar-toabsent20252026$\sim$ 2025-2026∼ 2025 - 2026) enables this study, with superior sensitivity to $Z^{\prime}$ around 10 MeV compared to CERN’s 160 GeV [106].
Quantum Entanglement: Quantum entanglement, a hallmark of quantum theory, has been observed in photons, leptons, and top quarks [107, 108, 109, 110, 111, 112]. In muon-electron and positron-electron scattering, entanglement is quantified via concurrence and CHSH inequality [92]. Simulations show strong CHSH violations at 1-10 GeV muon energies, with event rates of $\sim 2.6\times 10^{4}/day$ for 10 GeV muons and $\sim 1.9\times 10^{9}/s$ for 1 GeV positrons.
@@ IV.1 CP Violation
After the non-zeo $\theta_{13}$ was measured by Daya Bay in 2012 [118], the neutrino physics has entered a new era. This opens the possibility for experimental measurement of the Dirac CP phase $\delta_{D}$. The current measurements by T2K indicates that the leptonic CP is maximally violated wtih $\delta_{D}$ around $-90^{\circ}$ [119] for both mass orderings while NOvA prefers $180^{\circ}$ ($-90^{\circ}$) for the normal ordering (inverted ordering) [120]. However, the sensitivity of these two experiments is not high enough to fully determine the value of $\delta_{D}$. Especially, running both the neutrino and anti-neutrino modes at the first oscillation peak has several issues such as splitting the beam time, degeneracy between CP values, and intrinsically large CP uncertainty for the maximal CP phase [121, 122, 123, 124, 125, 126]. The next-generation experiments with optimization are desirable to establish the existence of leptonic CP violation and provide a precise measurement of $\delta_{D}$.
@@
If established, the HIAF-to-JUNO project would be the first accelerator neutrino experiment in China. With high-intensity neutrino beams and optimized neutrino energies, HIAF-to-JUNO would be sensitive to the Dirac CP phase. Especially for the maximal CP phases at ±$\pm 90^{\circ}$, its uncertainty tends to be largest inside the whole CP range for other accelerator experiments such as T2K/T2HK/NOvA/DUNE that are not optimized for the CP uncertainty. This leaves the HIAF-to-JUNO project a chance of optimizing its sensitivity for the maximal CP phase that is indicated by current measurements. This would allow HIAF-to-JUNO to become a very promising experiment for the precision CP measurement.
@@ IV.2 Coherent Scattering
The CEvNS cross section is the largest among neutrino scattering channels for neutrino energies below 100 MeV for most target nuclei. The COHERENT CsI(Na) detector accumulated 306$\pm$20 CEvNS candidate events with a 14.6 kg target mass, consistent with the Standard Model (SM) prediction of 341$\pm$11(theory)$\pm$42(experiment) [129]. However, this is insufficient for making efficient searches for new physics because of large statistical and systematic uncertainties.
@@
It expects to provide a complete picture of neutrino weak couplings predicted by the SM by making a precise measurement of the CEvNS cross sections at low momentum transfer. It allows probes of nuclear structure and improved constraints on the value of the weak nuclear charge. The CEvNS process is also useful for probing a yet unknown neutrino properties which can be only explained by the physics beyond the SM (BSM). They include searches for a neutrino magnetic moment, sterile neutrinos, and non-standard neutrino interactions (NSI) mediated by new particles. A noble CEvNS detector based on the dark-matter detection technique makes it plausible to observe extremely low-energy neutrinos which are undetectable because of their tiny cross-section. Due to the significantly enhanced rate for a lower energy neutrino, the CEvNS effect can be utilized to develop a small-size detector sensitive to neutrino emission in neutron stars and during stellar collapse. The CEvNS detector is also sensitive to solar neutrinos, atmospheric neutrinos, and Diffuse Supernova Neutrino Background which are irreducible backgrounds against direct searches for the dark matter.
High target masses of Cs and I atoms provide a substantially large coherence enhancement for a sufficient CEvNS event rate. Both target atom masses are similar and greatly simplify the detector response. The crystal exhibits a high light yield of $\sim$45 photons/keVee at 662 keVee, with a light emission peaking at 420 nm, The COHERENT experiment measured the light yield of 15.4$\pm$1.2 PE/keVee using the 241Am source [130] and achieved an energy threshold of 7 keVnr. An absolute energy scale was obtained as 1.2 PE/ keVnr [131] based on the measured quenching factor of $\sim$9%.
@@
The CICENNS detector expects to be located on the second floor of the experimental hall, approximately 25 m away from the spallation target. There exists a 10 m-concrete shielding structure around the beam target. The neutrino flux at the detector target is expected to be 5.4 ×$\times 10^{6}$ /cm2/s, roughly 11.4% of the flux at COHERENT (4.7 ×$\times 10^{7}$ /cm2/s at 20 m from the target for 1.4 MW) [131], and will become 5.4 ×$\times 10^{7}$ /cm2/s if the beam power is upgraded to 2.5 MW,
@@
A produced CEvNS event rate at CICENNS is shown in Fig. 6 as a function of nuclear recoil-energy threshold. The expected CEvNS production rate per 300 kg for a year is about 2500 (25000) events at the energy threshold of 7 keVnr and 3500 (35000) events at 5 keVnr, for a 250 kW (2.5 MW) beam. The detection efficiency is 60.5% above 5 keVnr while it rapidly increases from zero at 1.5 keVnr and becomes 35% at 3 keVnr. An expected CEvNS event rate at CiADS is 1520 (15200) events per 300 kg for a year data-taking at the energy threshold of 7 keVnr, for a 250 kW (2.5 MW) beam. The beam correlated background rate is estimated to be about 5% of a predicted CEvNS rate with an uncertainty of 25%. The steady-state background is estimated to be $4.1\times 10^{5}$ per year data-taking at energy threshold of 7 keVnr with its normalization uncertainty of 0.1%.
@@
The current uncertainty of the measured CEvNS cross section is +18 or -15% while that of the SM prediction is 4.8%. The 300 kg CsI(Na) CICENNS detector will reduce the uncertainty to roughly 5%, similar to the uncertainty of the nuclear form factor.
@@
The COHERENT obtained $\sin^{2}\theta_{\rm W}=0.220^{+0.028}_{-0.026}$ [129] compared to the SM prediction 0.23857(5). The current uncertainty from the CEvNS measurements is at the $\sim 10\%$ level at Q values of a few tens of MeV/c. The 300 kg CsI(Na) CICENNS detector will be able to make a percent-level ($\sim 1\%$) measurement.
@@
Based on the measured “flavored” CEvNS cross sections, it can provide a sensitive probe of BSM physics such as neutrino-quark vector NSIs, which can affect each neutrino flavor differently. Due to the sufficiently large CEvNS event rate, the CICENNS search results in significant improvement in the allowed parameter space.
@@
A neutrino magnetic moment less than $10^{-11}\mu_{B}$ can be measured by the modified the nuclear recoil energy distribution in response to CEvNS. However, an extremely low energy threshold below 1 keVnr and good energy resolution are required for the measurement.
@@ IV.3 Incoherent Scattering for Hadron physics
Furthermore, due to vacuum fluctuations, nucleons inherently contain $q\bar{q}$ components, but these are challenging to detect through conventional scattering processes. (Anti-)Neutrinos, by changing quark flavors, can convert $s\bar{s}$ components into ($s\bar{c}$)$c\bar{s}$. Thus, by identifying final states with net quark content of ($uuds\bar{c}$)$uudc\bar{s}$ in (anti)neutrino-proton reactions, it is possible to confirm their origin from (anti)neutrino scattering with ($\bar{s}$)$s$, thereby determining the $s\bar{s}$ component and its distribution in the proton. Similarly, the $u\bar{u}$ component can be identified through anti-neutrino scattering with final states of net quark content $uuud\bar{s}$, while the $d\bar{d}$ component can be determined via neutrino scattering with final states of net quark content $uudd\bar{c}$.
@@ IV.4 Three-dimensional PDFs
Neutrino deep inelastic scattering (DIS) is important in studying the nucleon and/or nucleus structures. It provides not only information on the flavor separation which cannot be realized in the charged lepton DIS experiments alone but also on electroweak physics by considering both the neutral-current and charged-current interactions. Measurable quantities in DIS are expressed in terms of parton distribution functions (PDFs) which reveal the longitudinal momentum distributions of partons or the one-dimensional structure of the nucleon. To explore the three-dimensional or the transverse momentum-dependent (TMD) PDFs, one is supposed to consider the semi-inclusive DIS (SIDIS) where a jet or a hadron is also measured in addition to the scattered lepton. Jet can be a direct probe of analyzing properties of the quark intrinsic transverse momentum by utilizing intrinsic asymmetries [133, 134]. Nevertheless, jet production SIDIS process can not cover the low energy kinematic regions or access the chiral-odd PDFs due to the conservation of the helicity [135], and has difficulties to identify light flavor PDFs even in charged-current scattering, since no hadron is measured or tagged. Therefore, one has to consider the hadron production SIDIS process where the cross section is given by the convolution of TMD PDFs and fragmentation functions (FFs). In this process, TMD PDFs and FFs are measured simultaneously with flavor identifications.
Neutrino DIS also play an important role in studying nuclear effects or nuclear (TMD) PDFs. In the inclusive process, the formalism of the neutrino nucleus scattering is the same to that of the neutrino nucleon scattering [136] and the target mass effect has been studied in depth [137]. However, systematic definitions of the nuclear TMD PDFs are still missing. Nuclear TMD PDFs are obtained by decomposing the nuclear two-point correlation function which is defined via nuclear state $|A\rangle$. Both of them need more researches.
@@ IV.5 Three-dimensional GPDs and TDAs of hadrons
where $M$ is a meson state, $W$ is the $W$ boson emitted from neutrino, and we use $N$ and $N_{1}$ to represent the nucleons in initial and final states, respectively. Since the cross section of neutrino DVCS is suppressed by an additional electromagnetic coupling constant $\alpha$ compared to the neutrino DVMP, thus, it will be more feasible to measure the latter by experiment [143, 144]. In these reactions, the energy of the neutrino beam should be larger than 5 GeV so that the QCD factorization condition ($-q^{2}\gg-t=-(p-p_{1})^{2}$) is satisfied, and the hadronic GPDs are extracted through the analysis of the measurements at the forward region (small-$t$ region).
Meanwhile, if we investigate the backward region (small-$u=(p-q)^{2}$ region) of these processes, and the nucleon-photon and nucleon-meson transition distribution amplitudes (TDA) can be also extracted from neutrino DVCS and DVMP processes, respectively [145]. Recently, the scientists at Jefferson Lab (JLab) just reported the first measurement of hadronic TDA using DVCS reaction with electron beams [146]. Similarly, one can investigate the TDAs of different quark flavors by measuring the neutrino DVCS and DVMP processes, which can be complementary to the ongoing measurements at JLab.
@@ V.1 Bound beta-decay of neutron measurement by microcalorimeter
The bound beta-decay (BoB) of neutron is also known as the two-body neutron decay, which is a rare decay mode into a hydrogen atom and an anti-neutrino. The state of neutrino can be exactly inferred by measuring the state of hydrogen atom, providing a possible pathway to explore new physics. However, this rare decay mode has not yet been observed so far, since it was predicted in 1947. The challenge in observing this decay is not only that its cross section is extremely low, equivalent to about branching ratio of the order of $10^{-6}$ of the three-body decay, but also that the final-state hydrogen atom is neutral and has extremely low kinetic energy, which cannot be effectively detected. Here we propose a microcalorimeter-based scheme for measuring the kinetic energies of hydrogen atoms produced from BoB of ultracold neutrons, which has a great advantage in terms of accuracy of the energy measurement.
For the BoB mode, the conservation of momentum and energy during the decay leads to the energy carried by the antineutrino is constant, while the kinetic energy of the hydrogen atom is also constant (325.7 eV),making it easy to be captured by solid materials. Since the hydrogen atom in the $n$s ($\geq 2$) state after the two-body decay can be captured by the detector, it will de-excite to the 1s state very quickly, transferring its thermal energy 10.2 eV more to the detector. Thus, the energy(kinetic energy plus potential energy) spectra of the hydrogen atoms on the detector would be a multiple line structure dominated by 325.7 eV and 335.9 eV. A microcalorimeter is a detector sensitive to heat release and has an extremely high energy resolution of the order of eV with no entrance window (dead layer) and offers the possibility of a wide range of available absorption materials [147]. The energy deposited by the particles is converted into heat and the resulting temperature rise is measured [148]. The high energy resolution of the order of eV allows to resolve the signal from emerging hydrogen atoms, which produce a narrow energy peak at 325.7 eV. Hydrogen atoms, electrons and protons can all produce heat signals in the detector. To get higher peak to background ratio, a 1kV anode will be used to remove Protons. A dilution refrigerator is needed by microcalorimeter.

### v2

@@ III.3 Probing and Knocking with Muons(PKMu)
The High Energy Fragment Separator beamline at HIAF will deliver GeV-scale muon beams [20]. As shown in Fig. 4, this enables studies on muon-philic DM [95, 96], charged lepton flavor violation(CLFV) [97], and quantum correlations in $\mu-e$ scattering [98, 99].
@@
PKMu experiment will employ Gas Electron Multiplier (GEM) [101] or Resistive Plate Chamber (RPC) [102] detectors for muon tracking. GEMs provide precise spatial resolution, widely applied in high-energy physics, including CMS [103, 104]. RPCs offer timing precision, robustness, and stability. Together, they offer a cost-effective platform for high-resolution tracking.
[…]
@@
This approach complements nuclear recoil searches(e.g., XENON1T, PandaX) [96, 107]. Detection efficiency remains key to improving limits. To further enhance the sensitivity of DM searches, we can conduct experiments untilizing muon beams with higher intensity and better focusing characteristic. It is crucial to emphasize that even a null result from this search would still be significant in placing new constraints on such scenarios. It would tighten the bounds on muon-philic DM interaction strengths.
[…]
Quantum Entanglement: Quantum entanglement, a hallmark of quantum theory, has been observed in photons, leptons, and top quarks [113, 114, 115, 116, 117, 118]. In muon-electron and positron-electron scattering, entanglement is quantified via concurrence and CHSH inequality [98]. Simulations show strong CHSH violations at 1-10 GeV muon energies, with event rates of $\sim 2.6\times 10^{4}/day$ for 10 GeV muons and $\sim 1.9\times 10^{9}/s$ for 1 GeV positrons. Detailed discussions can be found in [98].
@@ IV.1 CP Violation
After the non-zeo $\theta_{13}$ was measured by Daya Bay in 2012 [124], the neutrino physics has entered a new era. This opens the possibility for experimental measurement of the Dirac CP phase $\delta_{D}$. The current measurements by T2K indicates that the leptonic CP is maximally violated wtih $\delta_{D}$ around $90^{\circ}$ [125] for both mass orderings while NOvA prefers $180^{\circ}$ ($90^{\circ}$) for the normal ordering (inverted ordering) [126]. However, the sensitivity of these two experiments is not high enough to fully determine the value of $\delta_{D}$. Especially, running both the neutrino and anti-neutrino modes at the first oscillation peak has several issues such as splitting the beam time, degeneracy between CP values, and intrinsically large CP uncertainty for the maximal CP phase [127, 128, 129, 130, 131, 132]. The next-generation experiments with optimization are desirable to establish the existence of leptonic CP violation and provide a precise measurement of $\delta_{D}$.
@@
If established, the HIAF-to-JUNO project would be the first accelerator neutrino experiment in China. With high-intensity neutrino beams and optimized neutrino energies, HIAF-to-JUNO would be sensitive to the Dirac CP phase. Especially for the maximal CP phases at $\pm 90^{\circ}$, its uncertainty tends to be largest inside the whole CP range for other accelerator experiments such as T2K/T2HK/NOvA/DUNE that are not optimized for the CP uncertainty. This leaves the HIAF-to-JUNO project a chance of optimizing its sensitivity for the maximal CP phase that is indicated by current measurements. This would allow HIAF-to-JUNO to become a very promising experiment for the precision CP measurement.
@@ IV.2 Coherent Scattering
[…]
@@
It expects to provide a complete picture of neutrino weak couplings predicted by the SM by making a precise measurement of the CEvNS cross sections at low momentum transfer. It allows probes of nuclear structure and improved constraints on the value of the weak nuclear charge. The CEvNS process is also useful for probing a yet unknown neutrino properties which can be only explained by the physics beyond the SM (BSM). They include searches for a neutrino magnetic moment, sterile neutrinos, and non-standard neutrino interactions (NSI) mediated by new particles. A CEvNS detector based on the dark-matter detection technique makes it plausible to observe extremely low-energy neutrinos which are undetectable because of their tiny cross-section. Due to the significantly enhanced rate for a lower energy neutrino, the CEvNS effect can be utilized to develop a small-size detector sensitive to neutrino emission in neutron stars and during stellar collapse. The CEvNS detector is also sensitive to solar neutrinos, atmospheric neutrinos, and Diffuse Supernova Neutrino Background which are irreducible backgrounds against direct searches for the dark matter.
[…]
@@
The CICENNS detector expects to be located on the second floor of the experimental hall, approximately 25 m away from the spallation target. There exists a 10 m-concrete shielding structure around the beam target. The neutrino flux at the detector target is expected to be 5.4 $\times 10^{6}$ /cm2/s, roughly 11.4% of the flux at COHERENT (4.7 $\times 10^{7}$ /cm2/s at 20 m from the target for 1.4 MW) [137], and will become 5.4 $\times 10^{7}$ /cm2/s if the beam power is upgraded to 2.5 MW,
@@
A produced CEvNS event rate at CICENNS is shown in Fig. 6 as a function of nuclear recoil-energy threshold. The expected CEvNS production rate per 300 kg for a year is about 2500 (25000) events at the energy threshold of 7 keVnr and 3500 (35000) events at 5 keVnr, for a 250 kW (2.5 MW) beam. This would be the world-largest CEvNS data sample. The detection efficiency is 60.5% above 5 keVnr while it rapidly increases from zero at 1.5 keVnr and becomes 35% at 3 keVnr. An expected CEvNS event rate at CiADS is 1520 (15200) events per 300 kg for a year data-taking at the energy threshold of 7 keVnr, for a 250 kW (2.5 MW) beam. The beam correlated background rate is estimated to be about 5% of a predicted CEvNS rate with an uncertainty of 25%. The steady-state background is estimated to be $4.1\times 10^{5}$ per year data-taking at energy threshold of 7 keVnr with its normalization uncertainty of 0.1%.
@@
The current uncertainty of the measured CEvNS cross section is +18 or -15% while that of the SM prediction is 4.8%. The 300 kg CsI(Na) CICENNS detector will reduce the uncertainty to roughly 5%, similar to the uncertainty of the nuclear form factor. Figure 7 shows allowed contours in this parameter space, expected from the CICENNS and obtained from the COHERENT.
Figure 7: Contours for the flavored CEvNS cross section obtained by the CICENNS expectation (blue) in comparison with the COHERENT data (gray) [135]. An allowed area expected by the CICENNS is significantly reduced from the COHERENT.
@@
The COHERENT obtained $\sin^{2}\theta_{\rm W}=0.220^{+0.028}_{-0.026}$ [135] compared to the SM prediction 0.23857(5). The current uncertainty from the CEvNS measurements is at the $\sim 10\%$ level at Q values of a few tens of MeV/c. The 300 kg CsI(Na) CICENNS detector will be able to make a percent-level ($\sim 1\%$) measurement. Figure 8 shows measured weak mixing angles and the CICENNS expected sensitivity.
Figure 8: Comparison of measured weak mixing angles. The expected error from the CICENNS detector is roughly a percent level while it is $\sim$10% for the COHERENT.
@@
Figure 9: Extracted mean radius of neutron distribution (neutron radius) from the observed CEvNS nuclear recoil energy spectrum. The expected error from the CICENNS detector is roughly a percent level while it is $\sim$ 8% for the COHERENT.
@@
Based on the measured “flavored” CEvNS cross sections, it can provide a sensitive probe of BSM physics such as neutrino-quark vector NSIs, which can affect each neutrino flavor differently. The allowed contour in the $\epsilon^{u}_{ee}$ and $\epsilon^{u}_{\mu\mu}$ parameter space is shown in Fig. 10. Due to the sufficiently large CEvNS event rate, the CICENNS search results in significant improvement in the allowed parameter space.
[…]
@@
A neutrino magnetic moment less than $10^{-11}\mu_{B}$ can be measured by the modified the nuclear recoil energy distribution in response to CEvNS. However, an extremely low energy threshold below 1 keVnr and good energy resolution are required for the measurement. The measurement sensitivity of CICENNS in comparison with COHERENT is shown in Fig. 11.
Figure 11: Comparison of CICENNS and COHERENT sensitivities on search for neutrino magnetic moment. Left: sensitivity on $\nu_{e}$ magnetic moment, Right: sensitivity on $\nu_{\mu}$ magnetic moment.
@@ IV.3 Incoherent Scattering for Hadron physics
[…]
@@ IV.4 Three-dimensional PDFs
The energy of HIAF neutrinos can reach 10 GeV, which is suitable for deep inelastic scattering (DIS) measurements. The DIS of neutrino plays an important role in studying partonic structures of nucleons or nuclei. It provides not only information on the flavor separation which cannot be realized in the charged lepton DIS experiments alone but also on electroweak physics by considering both the neutral-current and charged-current interactions. Measurable quantities in DIS are expressed in terms of parton distribution functions (PDFs) which reveal the longitudinal momentum distributions of partons or the one-dimensional structure of the nucleon. To explore the three-dimensional or the transverse momentum-dependent (TMD) PDFs, one is supposed to consider the semi-inclusive DIS (SIDIS) where a jet or a hadron is also measured in addition to the scattered lepton. Jet can be a direct probe of analyzing properties of the quark intrinsic transverse momentum by utilizing intrinsic asymmetries [139, 140]. Nevertheless, jet production SIDIS process can not cover the low energy kinematic regions or access the chiral-odd PDFs due to the conservation of the helicity [141], and has difficulties to identify light flavor PDFs even in charged-current scattering, since no hadron is measured or tagged. It is therefore to consider the hadron production SIDIS process where the cross section is given by the convolution of TMD PDFs and fragmentation functions (FFs). In this process, TMD PDFs and FFs are measured simultaneously with flavor identifications. Specially, measuring the dimuon signals in the charged-current SIDIS can provide promising direct experimental constraints on the strange and anti-strange quark distributions of the nucleon [142, 143].
Neutrino DIS also play an important role in studying nuclear effects or nuclear TMD PDFs. In the inclusive process, the formalism of the neutrino nucleus scattering is the same to that of the neutrino nucleon scattering [144] and the target mass effect has been studied in depth [145]. However, systematic definitions of the nuclear TMD PDFs are still missing. Nuclear TMD PDFs are obtained by decomposing the nuclear two-point correlation function which is defined via nuclear state $|A\rangle$. Both of them need more researches.
@@ IV.5 Three-dimensional GPDs and TDAs of hadrons
where $M$ is a meson state, $W$ is the $W$ boson emitted from neutrino, and we use $N$ and $N_{1}$ to represent the nucleons in initial and final states, respectively. Since the cross section of neutrino DVCS is suppressed by an additional electromagnetic coupling constant $\alpha$ compared to the neutrino DVMP, thus, it will be more feasible to measure the latter by experiment [151, 152]. In these reactions, the energy of the neutrino beam should be larger than 5 GeV so that the QCD factorization condition ($q^{2}\gg-t=-(p-p_{1})^{2}$) is satisfied, and the hadronic GPDs are extracted through the analysis of the measurements at the forward region (small-$t$ region).
[…]
@@ V.1 Bound beta-decay of neutron measurement by microcalorimeter
The bound beta-decay (BoB) of neutron is a rare decay mode into a hydrogen atom and an anti-neutrino. The state of neutrino can be exactly inferred by measuring the state of hydrogen atom, providing a possible pathway to explore new physics. However, this rare decay mode has not yet been observed so far, since it was predicted in 1947. The challenge in observing this decay is not only that its cross section is extremely low, equivalent to about branching ratio of the order of $10^{-6}$ of the three-body decay, but also that the final-state hydrogen atom is neutral and has extremely low kinetic energy, which cannot be effectively detected. Here we propose a microcalorimeter-based scheme for measuring the kinetic energies of hydrogen atoms produced from BoB of ultracold neutrons, which has a great advantage in terms of accuracy of the energy measurement.
For the BoB mode, the conservation of momentum and energy during the decay leads to the energy carried by the antineutrino is constant, while the kinetic energy of the hydrogen atom is also constant (325.7 eV),making it easy to be identify. Since the hydrogen atom in the $n$s ($\geq 2$) state after the two-body decay can be captured by the detector, it will de-excite to the 1s state very quickly, transferring its thermal energy 10.2 eV more to the detector. Thus, the energy(kinetic energy plus potential energy) spectra of the hydrogen atoms on the detector would be a multiple line structure dominated by 325.7 eV and 335.9 eV. A microcalorimeter is a detector sensitive to heat release and has an extremely high energy resolution of the order of eV with no entrance window (dead layer) and offers the possibility of a wide range of available absorption materials [155]. The high energy resolution of the order of eV allows to resolve the signal from emerging hydrogen atoms, which produce a narrow energy peak at 325.7 eV. Hydrogen atoms, electrons and protons can all produce heat signals in the detector. To get higher peak to background ratio, a 1kV anode will be used to remove Protons. The 2s H atoms will de-excited to 1s in strong electric field, only 1s peak is enough for aim to observe BOB, We need find a way to remove Protons without de-excite 2s H.